# On circumcentered direct methods for monotone variational inequality problems

### v1

@@
Abstract
The variational inequality problem (VIP) plays a central role in the theory and applications in continuous optimization. In particular, minimization problems and KKT systems can be regard as VIPs. In this work, we present the first methods using circumcenter iterations for solving VIPs. The circumcentered-reflection method (CRM) is a tool based on projections developed with the aim of finding a point in the intersection of finitely many closed convex sets. CRM has gone through enhancements and adaptations over the last few years and was proven to be faster in many settings than competitors such as alternating projections and the Douglas-Rachford method. One of the nice features of CRM is that it is able to deal with approximate projections, which is exactly what we enforced in this article theoretically and numerically. We present both a circumcenter method for the VIP with paramonotone operator and monotone operator, the first employing exact projections and the second approximate ones. Convergence results showing their convergence are established. Numerically, our experiments indicate good performance, including advantages over the well-know extragradient method.
1 Introduction
1.1 Variational inequalities
[…]
$\langle F(x^{*}),x-x^{*}\rangle\geq 0,~{}~{}\forall x\in C.$ (1.1)
The set of solutions of VIP$(F,C)$ will be denoted by SOL$(F,C)$. In particular, if $F=\nabla g$ for a convex function $g:\mathds{R}^{n}\to\mathds{R}$, then VIP$(F,C)$ is equivalent to the problem of minimizing $g$ over $C$.
@@ 1.1 Variational inequalities
The variational inequality problem was initially introduced and studied by [36]. Variational inequalities have attracted wide attention, since they cover a large variety of problems arising in different fields, such as economics, structural analysis, optimization, engineering sciences, and operations research, see [2, 20, 22, 34, 35, 44, 36, 51] and the references therein. An excellent survey of methods for variational inequality problems can be found in [30].
1.2 Gradient-type methods for variational inequalities
@@
Here, we are interested in direct methods for solving VIP($F,C$). The basic idea consists of extending the projected gradient method for constrained optimization, i.e., for the problem of minimizing $g(x)$ subject to $x\in C$. This problem is a particular case of VIP($F,C$) with $F=\nabla g$. This procedure is given by the following iterative scheme:
$x^{0}\in C,$ (1.2)
$x^{k+1}=P_{C}(x^{k}-\alpha_{k}\nabla g(x^{k})),$ (1.3)
@@
with $\alpha_{k}>0$ for all $k$. The coefficients $\alpha_{k}$ are called stepsizes and $P_{C}:\mathds{R}^{n}\to C$ is the orthogonal projection onto $C$, i.e., $P_{C}(x)={\rm argmin}_{y\in C}\left\lVert x-y\right\rVert$.
An immediate extension of the method (1.2)–(1.3) to VI$(F,C)$ is the iterative procedure given by
$x^{0}\in C,$ (1.4)
$x^{k+1}=P_{C}(x^{k}-\alpha_{k}F(x^{k})),$ (1.5)
which will be called Gradient-Type (GT) method for variational inequalities.
@@
Convergence results for this method require some monotonicity properties of $F$. We introduce below several possible options.
@@
Consider an operator $F:\mathds{R}^{n}\rightarrow\mathds{R}^{n}$ and a convex set $M\subset\mathds{R}^{n}$. $F$ is said to be
  • (i)
    monotone on $M$ if $\langle F(x)-F(y),x-y\rangle\geq 0$ for all $x,y\in M$;
  • (ii)
    paramonotone on $M$ if it is monotone in $M$, and whenever $\langle x-y,F(x)-F(y)\rangle=0$ with $x,y\in M$, it holds that $F(x)=F(y)$;
  • (iii)
    strictly monotone on $M$ if $\langle x-y,F(x)-F(y)\rangle>0$ for all $x,y\in M$, $x\neq y$;
  • (iv)
    uniformly monotone on $M$ if $\langle x-y,F(x)-F(y)\rangle\geq\varphi(\left\lVert F(x)-F(y)\right\rVert)$ for all $x,y\in M$, where $\varphi:\mathds{R}_{+}\to\mathds{R}$ is an increasing function with $\varphi(0)=0$;
  • (v)
    strongly monotone on $M$ if $\langle x-y,F(x)-F(y)\rangle\geq\omega\left\lVert F(x)-F(y)\right\rVert^{2}$ for all $x,y\in M$ and some $\omega>0$.
Remark: It follows from Definition 1.1 that the following implications hold: (iv) $\Rightarrow$ (iii) $\Rightarrow$ (ii) $\Rightarrow$ (i). The reverse assertions are not true in general.
We mention that all these monotonicity notions can be defined also for point-to-set operators, i.e., for $F:\mathds{R}^{n}\to{\mathcal{P}}(\mathds{R}^{n})$. In this situation, $F$ is usually required to be maximal monotone. In the point-to-point setting (the one considered in this paper), maximal monotonicity reduces to monotonicity and continuity. We make this observation because the proofs of some results that will be invoked later require maximal monotonicity of $F$. Our continuity and monotonicity assumptions ensure that these results hold in our case.
It has been proved in [31] that when $F$ is strongly monotone and Lipschitz continuous, i.e., there exists $L>0$ such that $\left\lVert F(x)-F(y)\right\rVert\leq L\left\lVert x-y\right\rVert$ for all $x,y\in\mathds{R}^{n}$, then the GT method converges to the unique solution of VIP$(F,C)$, provided that $\alpha_{k}\in\left(\epsilon,\frac{2\omega}{L^{2}}\right)$ for all $k$ and for some $\epsilon>0$.
These results were later improved under weaker monotonicity assumptions on the operator $F$. The case of uniformly monotone operators was analyzed in [1] and the case of paramonotone ones in [18]. The first algorithm proposed in this paper is shown to converge when $F$ is paramonotone. Thus, it is worthwhile to look more carefully at this class of operators.
@@
    If $F:\mathds{R}^{n}\to\mathds{R}^{n}$ is monotone and differentiable, and $J_{F}(x)$ denotes the Jacobian matrix of $F$ at $x$, then $F$ is paramonotone if and only if ${\rm Rank}(J_{F}(x)+J_{F}(x)^{t})={\rm Rank}(J_{F}(x))$ for all $x$, see Proposition $4.2$ in [39].
@@
It follows that affine operators of the form $F(x)=Ax+b$ are paramonotone when $A$ is positive semidefinite (not necessarily symmetric), and ${\rm Rank}(A+A^{t})={\rm Rank}(A)$. This situation includes cases of nonsymmetric and singular matrices, where the solution set of VI$(F,C)$ can be a subspace, differently from the case of strictly or strongly monotone operators, for which the solution set is always a singleton, when nonempty. Of course, this can also happen for nonlinear operators, e.g., when $F=\nabla g$ for a convex $g$, in which case the solution set is the set of minimizers of $g$ on $C$, which need not be a singleton. On the other hand, when $A$ is nonzero and skew-symmetric, $F(x)=Ax+b$ is monotone but not paramonotone.
[…]
@@
In order to overcome this difficulty, the following iteration, called extragradient method, was proposed by G.M. Korpelevich in [45]:
$y^{k}=P_{C}\left(x^{k}-\beta F(x^{k})\right),$ (1.6)
$x^{k+1}=P_{C}\left(x^{k}-\beta F(y^{k})\right),$ (1.7)
@@
where $\beta>0$ is a fixed number. Assuming that $F$ is monotone and Lipschitz continuous with constant $L$, that $\beta\in\left(\epsilon,\frac{1}{L}\right)$ for some $\epsilon>0$, and that VIP($F,C$) has solutions, Korpelevich showed that the sequence $\{x^{k}\}$ generated by (1.6)-(1.7) converges to a solution of VIP$(F,C)$.
Our second algorithm, which converges under the sole assumption that $F$ is monotone, is based on this extragradient method.
1.3 The case of $C=\cap_{i=1}^{m}C_{i}$
@@
In principle, these methods are direct, in the sense there is no subproblem to be solved at each iteration, which requires just evaluation of $F$ and computation of the projection $P_{C}$ at certain points. However, the computation of the projection onto a convex set, in general, is not elementary (except for some very special convex sets, such as balls or halfspaces), and hence it is worthwhile to explore situations in which the projection step can be replaced by less expensive procedures. A typical situation is when the convex set $C$ is given as intersections of simpler sets, i.e., $C=\cap_{i=1}^{m}C_{i}$, where each $C_{i}\subset\mathds{R}^{n}$ is closed and convex. For instance, in the very frequent case when $C=\{x\in\mathds{R}^{n}:g_{i}(x)\leq 0\,(1\leq i\leq m)\}$, with convex $g_{i}:\mathds{R}^{n}\to\mathds{R}$, it is natural to take $C_{i}=\{x\in\mathds{R}^{n}:g_{i}(x)\leq 0\}$. Usually, $P_{C_{i}}$ will be much easier to compute than $P_{C}$.
@@
A relaxed projection algorithm proposed by He and Yang in [37] for solving VI$(F,C)$ is as follows:
$x^{k+1}=\left(P_{m}^{k}\circ\cdots\circ P_{2}^{k}\circ P_{1}^{k}\right)(x^{k}-\alpha_{k}F(x^{k})).$ (1.8)
@@
  • i)
    We consider a GT method which converges for paramonotone operators (rather than strongly monotone ones), under a standard coerciveness assumption, and an extragradient-type method which converges for plainly monotone operators
  • ii)
    We use simultaneous (rather than sequential) projections onto the sets $C_{i}$.
  • iii)
@@
1.4 Circumcentered-Reflection methods
The Convex Feasibility Problem (CFP) consists of finding a point in the intersection of $m$ convex subsets of $\mathds{R}^{n}$.
Two classical methods for CFP are the Sequential and the Simultaneous Projection Methods (SePM and SiPM respectively, which originate in [42] and [24]).
@@ 1.4 Circumcentered-Reflection methods
$x^{k+1}=\left(P_{m}\circ\cdots\circ P_{2}\circ P_{1}\right)(x^{k}),$ (1.9)
@@
$x^{k+1}:=T_{A,B}(x^{k})={\rm circ}(x^{k},R_{A}(x^{k}),R_{B}(R_{A}(x^{k}))),$ (1.11)
where $P_{A}$, $P_{B}$ are the projection operators onto $A,B$, $R_{A}=2P_{A}-{\rm I}$ and $R_{B}=2P_{B}-{\rm I}$ are the standard reflection operators, ${\rm circ}(x^{k},R_{A}(x^{k}),R_{B}(R_{A}(x^{k})))$ is the circumcenter of $x^{k}$, $R_{A}(x^{k})$ and $R_{B}(R_{A}(x^{k}))$, i.e., a point equidistant to these three points and laying in the affine manifold spanned by them. It has been proved in [14] that the sequence generated in this way converges to a point in $C$ if $C$ is nonempty and either $A$ or $B$ is an affine manifold.
@@
${\textbf{W}}:=C_{1}\times\cdots\times C_{m}\subset\mathds{R}^{nm},$ (1.12)
@@
${\textbf{D}}:=\{{\textbf{y}}\in\mathds{R}^{nm}:\textbf{y}=(x,\ldots,x),x\in\mathds{R}^{n}\}.$ (1.13)
@@
$(P_{\textbf{D}}\circ P_{\textbf{W}})({\textbf{y}})=(\bar{z},\dots,\bar{z})$
[…]
${\rm\mathbf{T}_{\textbf{W},\textbf{D}}}({\textbf{z}})={\rm circ}(\mathbf{R}_{\mathbf{D}}\mathbf{R}_{\mathbf{W}}(\mathbf{z}),\mathbf{R}_{\mathbf{W}}(\mathbf{z}),\mathbf{z}),$ (1.14)
@@
$\mathbf{R}_{\mathbf{W}}(\mathbf{z})=\left(R_{C_{1}}(z),\cdots,R_{C_{m}}(z)\right)$ (1.15)
@@
$\mathbf{R}_{\mathbf{D}}(\mathbf{z})=\newline \left(2\left(\sum_{i=1}^{m}P_{C_{i}}(z)\right)-z,\ldots,2\left(\sum_{i=1}^{m}P_{C_{i}}(z)\right)-z\right),$ (1.16)
with ${\mathbf{z}}=(z,\dots,z)\in{\mathbf{D}}$. In principle, the operator $\mathbf{T}_{\mathbf{W},\mathbf{D}}$ is defined from $\mathds{R}^{nm}$ to $\mathds{R}^{mn}$. However by Proposition 2.7(i) in [5], we know that if $\mathbf{z}\in\mathbf{D}$, then $\mathbf{T}_{\mathbf{W},\mathbf{D}}(\mathbf{z})\in\mathbf{D}$. Therefore, we conclude that $\mathbf{T}_{\mathbf{W},\mathbf{D}}:\mathbf{D}\rightarrow\mathbf{D}.$
The CRM method has proved to be quite superior to classical projection methods for solving the CRM, achieving superlinear convergence under mild assumptions, while the classical methods are at most linearly convergent; see [5, 4, 10, 11, 12, 13, 6, 7, 9, 8, 27, 46, 48, 47].
1.5 Approximate projections
CRM is superior to most known methods for CFP, but still has a serious drawback: the computation of the projections onto the sets $C_{i}$. A variant of CRM which uses approximate projections in the product space setting was introduced in [3] and denoted as ACRM (Approximate CRM). We begin by defining separating sets for convex sets.

### v2

@@
Abstract.
Circumcentered techniques have been shown to significantly accelerate projection-based methods for convex feasibility problems. Motivated by this success, we propose two direct methods with circumcenter acceleration for solving variational inequality problems involving two classes of operators: paramonotone and monotone. Both schemes rely on approximate projections onto separating halfspaces, thereby avoiding computationally expensive exact projections. We establish convergence results for both methods and conduct numerical experiments, demonstrating that the proposed algorithms outperform classical methods, such as the extragradient algorithm, by orders of magnitude in terms of computational time, particularly when the feasible set is a complex intersection of convex sets.
Key words and phrases:
Variational inequalities, circumcentered-reflection method, monotone operators, paramonotone operators, projection methods
2020 Mathematics Subject Classification:
Primary 65K15, 90C33; Secondary 49J40, 47H05
R. Behling was partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico – CNPq (Grants XXXX 407147/2023-3 and 403197/2025-2) and Fundação de Amparo à Pesquisa do Estado do Rio de Janeiro – FAPERJ (Grant E-26/201.345/2021).
Y. Bello-Cruz was partially supported by the National Science Foundation (Grant DMS-2307328).
D. Liu was partially supported by CNPq (Grant 153172/2024-0).
L.-R. Santos was partially supported by CNPq (Grants 310571/2023-5, 407147/2023-3 and 403197/2025-2) and by the Fundação de Amparo à Pesquisa e Inovação do Estado de Santa Catarina – FAPESC (Edital 21/2024, Grant 2024TR002238).
1. Introduction
1.1. Variational inequalities
Let $C\subset\mathds{R}^{n}$ be a nonempty, closed, and convex set, and let $F:\mathds{R}^{n}\to\mathds{R}^{n}$ be a continuous operator. The variational inequality problem $\operatorname{VIP}(F,C)$ consists of finding some $x^{*}\in C$ such that
(1.1) $\langle F(x^{*}),x-x^{*}\rangle\geq 0,~~\forall x\in C.$
The set of solutions of $\operatorname{VIP}(F,C)$ will be denoted by $\operatorname{SOL}(F,C)$. In particular, if $F$ is the gradient operator $\nabla g$ of a convex and differentiable function $g:\mathds{R}^{n}\to\mathds{R}$, then $\operatorname{VIP}(F,C)$ reduces to the first-order optimality condition for minimizing $g$ over $C$. Here and henceforth, we denote by $\langle\cdot,\cdot\rangle$ and $\left\lVert\cdot\right\rVert$ the usual Euclidean inner product and norm, respectively.
@@ 1.1. Variational inequalities
The variational inequality problem was initially introduced and studied by [38]. Since then, this problem has attracted considerable attention, as it covers a wide range of problems arising in different fields. For instance, many models in economics, structural analysis, optimization, engineering sciences, and operations research can be posed as a VIP; see [2, 20, 22, 35, 37, 38, 45, 36, 41] and the references therein. An excellent survey of methods for VIPs can be found in [32].
1.2. Gradient-Type methods for variational inequalities
@@
In this work, we are interested in direct methods for solving $\operatorname{VIP}(F,C)$. The basic idea consists of extending the projected gradient method for constrained optimization, i.e., for the problem of minimizing $g(x)$ subject to $x\in C$. This problem is a particular case of, $\operatorname{VIP}(F,C)$ with $F=\nabla g$. This procedure is given by the following iterative scheme:
(1.2) $\begin{cases}x^{0}\in C,\\ x^{k+1}=P_{C}(x^{k}-\alpha_{k}\nabla g(x^{k})),\end{cases}$
@@
with $\alpha_{k}>0$ for all $k$. The coefficients $\alpha_{k}$ are called stepsizes and $P_{C}:\mathds{R}^{n}\to C$ is the orthogonal projection onto $C$, i.e., $P_{C}(x)=\operatorname*{arg\,min}_{y\in C}\left\lVert x-y\right\rVert$.
An immediate extension of the method (1.2) to $\operatorname{VIP}(F,C)$ is the iterative procedure given by
(1.3) $\begin{cases}x^{0}\in C,\\ x^{k+1}=P_{C}(x^{k}-\alpha_{k}F(x^{k})),\end{cases}$
@@
which will be called Gradient-Type (GT) method for variational inequalities. Convergence results for this method require some monotonicity properties of $F$. We introduce below several possible options.
@@
Consider an operator $F:\mathds{R}^{n}\rightarrow\mathds{R}^{n}$ and a convex set $C\subset\mathds{R}^{n}$. $F$ is said to be
  1. (i)
    monotone on $C$ if $\langle F(x)-F(y),x-y\rangle\geq 0$ for all $x,y\in C$;
  2. (ii)
    paramonotone on $C$ if it is monotone on $C$, and whenever $\langle x-y,F(x)-F(y)\rangle=0$ with $x,y\in C$, it holds that $F(x)=F(y)$;
  3. (iii)
    strictly monotone on $C$ if $\langle x-y,F(x)-F(y)\rangle>0$ for all $\in C, xyx\neq y$;
  4. (iv)
    uniformly monotone on $C$ if $\langle x-y,F(x)-F(y)\rangle\geq\varphi(\left\lVert x-y\right\rVert)$ for all $x,y\in C$, where $\varphi:\mathds{R}_{+}\to\mathds{R}$ is an increasing function with $\varphi(0)=0$;
  5. (v)
    strongly monotone on $C$ if $\langle x-y,F(x)-F(y)\rangle\geq\omega\left\lVert x-y\right\rVert^{2}$ for all $x,y\in C$ and some $\omega>0$.
Remark 1.2.
It follows from Definition 1.1 that the following implications hold: (iv) $\Rightarrow$ (iii) $\Rightarrow$ (ii) $\Rightarrow$ (i). The reverse assertions are not true in general.
We mention that all these monotonicity notions can also be defined also for point-to-set operators [27], i.e., for $F:\mathds{R}^{n}\to{\mathcal{P}}(\mathds{R}^{n})$. In this situation, $F$ is usually required to be maximal monotone [7, Definition 20.20]. In the point-to-point setting (the one considered in this paper), maximal monotonicity reduces to monotonicity and continuity. We make this observation because the proofs of some results that will be invoked later require maximal monotonicity of $F$. Our continuity and monotonicity assumptions ensure that those results remain valid in our case.
It has been proved in [54] that when $F$ is strongly monotone and Lipschitz continuous, i.e., there exists $L>0$ such that $\left\lVert F(x)-F(y)\right\rVert\leq L\left\lVert x-y\right\rVert$ for all $x,y\in\mathds{R}^{n}$, then the GT method converges to the unique solution of $\operatorname{VIP}(F,C)$, provided that $\alpha_{k}\in\left(\epsilon,\frac{2\omega}{L^{2}}\right)$ for all $k$ and for some $\epsilon>0$, where $\omega>0$ is the strong monotonicity constant.
These results were later improved under weaker monotonicity assumptions on the operator $F$. The case of uniformly monotone operators was analyzed in [1] and the case of paramonotone operators in [18]. The first algorithm proposed in this paper is shown to converge under the assumption that $F$ is paramonotone. Thus, it is worthwhile to look more carefully at this class of operators.
@@
    If $F:\mathds{R}^{n}\to\mathds{R}^{n}$ is monotone and differentiable, and $J_{F}(x)$ denotes the Jacobian matrix of $F$ at $x$, then $F$ is paramonotone if and only if $\operatorname{rank}(J_{F}(x)+J_{F}(x)^{\top})=\operatorname{rank}(J_{F}(x))$ for all $x$; see Proposition $4.2$ in [42].
@@
It follows that affine operators of the form $F(x)=Ax+b$ are paramonotone when $A$ is positive semidefinite (not necessarily symmetric), and $\operatorname{rank}(A+A^{\top})=\operatorname{rank}(A)$. This situation includes cases of nonsymmetric and singular matrices, where the solution set of $\operatorname{VIP}(F,C)$ can be a subspace, in contrast to the case of strictly or strongly monotone operators, for which the solution set is always a singleton, when nonempty [32]. Of course, this can also happen for nonlinear operators, e.g., when $F=\nabla g$ for a convex function $g$, in which case the solution set is the set of minimizers of $g$ on $C$, which need not be a singleton. On the other hand, when $A$ is nonzero and skew-symmetric, $F(x)=Ax+b$ is monotone but not paramonotone.
Remark 1.3.
There is no general way to relax the assumption on $F$ to plain monotonicity. For example, consider $F:\mathds{R}^{2}\to\mathds{R}^{2}$ defined as $F(x)=Ax$, with $A=\left(\begin{array}[]{rr}0&\,\,1\\ -1&\,\,0\\ \end{array}\right)$. $F$ is monotone, and the unique solution of $\operatorname{VIP}(F,C)$ is $x^{*}=0$. However, it is easy to check that $\left\lVert x^{k}-\alpha_{k}F(x^{k})\right\rVert>\left\lVert x^{k}\right\rVert$ for all $x^{k}\neq 0$ and all $\alpha_{k}>0$, and therefore the sequence generated by the GT method moves away from the solution, regardless of the choice of the stepsize $\alpha_{k}$.
@@
In order to overcome this difficulty, the following iteration, called the extragradient method, was introduced by G.M. Korpelevich in [46]:
(1.4) $\begin{cases}y^{k}=P_{C}\left(x^{k}-\beta F(x^{k})\right),\\ x^{k+1}=P_{C}\left(x^{k}-\beta F(y^{k})\right),\end{cases}$
@@
where $\beta>0$ is a fixed number. Assuming that $F$ is monotone and Lipschitz continuous with constant $L$, that $\beta\in\left(\epsilon,\frac{1}{L}\right)$ for some $\epsilon>0$, and that $\operatorname{VIP}(F,C)$ has solutions, Korpelevich showed that the sequence $\{x^{k}\}$ generated by (1.4) converges to a solution of $\operatorname{VIP}(F,C)$. Numerous variants and improvements of this method have been proposed in the literature; see, for instance, [53, 48, 23, 55, 40].
1.3. The case of $C=\cap_{i=1}^{m}C_{i}$
@@
In principle, these methods are direct, in the sense that there is no subproblem to be solved at each iteration, which requires just the evaluation of $F$ and the computation of the projection $P_{C}$ at certain points. However, the computation of the projection onto a convex set, in general, is not elementary (except for some very special convex sets, such as balls or halfspaces), and hence it is worthwhile to explore situations in which the projection step can be replaced by less expensive procedures. A typical situation is when the convex set $C$ is given as an intersection of simpler sets, i.e., $C=\cap_{i=1}^{m}C_{i}$, where each $C_{i}\subset\mathds{R}^{n}$ is closed and convex. For instance, in the very frequent case when $C=\{x\in\mathds{R}^{n}:g_{i}(x)\leq 0\,(1\leq i\leq m)\}$, with convex $g_{i}:\mathds{R}^{n}\to\mathds{R}$, it is natural to take $C_{i}=\{x\in\mathds{R}^{n}:g_{i}(x)\leq 0\}$. Usually, $P_{C_{i}}$ will be much easier to compute than $P_{C}$.
@@
A relaxed projection algorithm proposed by He and Yang in [39] for solving $\operatorname{VIP}(F,C)$ is as follows:
(1.5) $x^{k+1}=\left(P_{m}^{k}\circ\cdots\circ P_{2}^{k}\circ P_{1}^{k}\right)(x^{k}-\alpha_{k}F(x^{k})),$
@@
  1. (i)
    We consider a GT method which converges for paramonotone operators (rather than strongly monotone ones), under a standard coerciveness assumption, and another method which converges for plainly monotone operators.
  2. (ii)
    We use simultaneous (rather than sequential) projections onto separating sets (e.g. halfspaces) of the sets $C_{i}$.
  3. (iii)
@@
1.4. Circumcentered-Reflection methods
The Convex Feasibility Problem (CFP) consists of finding a point in the intersection of $m$ convex subsets of $\mathds{R}^{n}$. Two classical methods for CFP are the Sequential and the Simultaneous Projection Methods (SePM and SiPM, respectively), which originate in [44] and [26].
@@ 1.4. Circumcentered-Reflection methods
(1.6) $x^{k+1}=\left(P_{m}\circ\cdots\circ P_{2}\circ P_{1}\right)(x^{k}),$
@@
(1.8) $x^{k+1}:=T_{A,B}(x^{k})=\operatorname*{circ}(x^{k},R_{A}(x^{k}),R_{B}(R_{A}(x^{k}))),$
where $P_{A}$, $P_{B}$ are the projection operators onto $A,B$, $R_{A}=2P_{A}-{\rm I}$ and $R_{B}=2P_{B}-{\rm I}$ are the standard reflection operators. The point $\operatorname*{circ}(x^{k},R_{A}(x^{k}),R_{B}(R_{A}(x^{k})))$ is the circumcenter of $x^{k}$, $R_{A}(x^{k})$ and $R_{B}(R_{A}(x^{k}))$, i.e., a point equidistant from these three points and lying in the affine manifold spanned by them. One can regard $\operatorname*{circ}(\cdot,\cdot,\cdot)$ as a mapping from $\mathds{R}^{n}\times\mathds{R}^{n}\times\mathds{R}^{n}$ to $\mathds{R}^{n}$. It has been proved in [14] that the sequence generated in this way converges to a point in $C$ if $C$ is nonempty and either $A$ or $B$ is an affine manifold.
@@
(1.9) ${\textbf{W}}\coloneqq C_{1}\times\cdots\times C_{m}\subset\mathds{R}^{nm},$
@@
(1.10) ${\textbf{D}}\coloneqq\{{\textbf{y}}\in\mathds{R}^{nm}:\textbf{y}=(x,\ldots,x),x\in\mathds{R}^{n}\}.$
@@
$(P_{\textbf{D}}\circ P_{\textbf{W}})({\textbf{y}})=(\bar{z},\dots,\bar{z}),$
with $\bar{z}=\frac{1}{m}\sum_{i=1}^{m}P_{i}(z)$. Noting that if we consider the SiPM iteration with $x^{k}=z$ we get $x^{k+1}=\bar{z}$, we conclude that an iteration of SePM in the product space $\mathds{R}^{nm}$ with the two convex sets ${\textbf{W}},{\textbf{D}}$, starting at a point in D, is equivalent to an iteration of SiPM in $\mathds{R}^{n}$ with the sets $C_{1},\dots,C_{m}$. Note also that D is an affine manifold (indeed, a linear subspace). Hence, applying CRM with the sets ${\textbf{W}},{\textbf{D}}\subset\mathds{R}^{nm}$ instead of $A,B\subset\mathds{R}^{n}$, we get a sequence ${\textbf{z}}^{k}=(x^{k},\dots,x^{k})\in\mathds{R}^{nm}$, with $\{x^{k}\}\subset\mathds{R}^{nm}$ converging to a point $x^{*}\in C=\cap_{i=1}^{m}C_{i}$, where the $C_{i}$’s are now arbitrary closed and convex subsets of $\mathds{R}^{n}$. We define now the circumcentered-reflection operator in the product space as follows:
(1.11) ${\rm\mathbf{T}_{\textbf{W},\textbf{D}}}({\textbf{z}})=\operatorname*{circ}(\mathbf{R}_{\mathbf{D}}(\mathbf{R}_{\mathbf{W}}(\mathbf{z})),\mathbf{R}_{\mathbf{W}}(\mathbf{z}),\mathbf{z}),$
@@
(1.12) $\mathbf{R}_{\mathbf{W}}(\mathbf{z})=\left(R_{C_{1}}(z),\ldots,R_{C_{m}}(z)\right),$
@@
(1.13) $\mathbf{R}_{\mathbf{D}}(\mathbf{z})=\newline \left(\frac{2}{m}\left(\sum_{i=1}^{m}P_{C_{i}}(z)\right)-z,\ldots,\frac{2}{m}\left(\sum_{i=1}^{m}P_{C_{i}}(z)\right)-z\right),$
with ${\mathbf{z}}=(z,\dots,z)\in{\mathbf{D}}$. In principle, the operator $\mathbf{T}_{\mathbf{W},\mathbf{D}}$ is defined from $\mathds{R}^{nm}$ to $\mathds{R}^{mn}$. However, by Proposition 2.7(i) in [4], we know that if $\mathbf{z}\in\mathbf{D}$, then $\mathbf{T}_{\mathbf{W},\mathbf{D}}(\mathbf{z})\in\mathbf{D}$. Therefore, we conclude that $\mathbf{T}_{\mathbf{W},\mathbf{D}}:\mathbf{D}\rightarrow\mathbf{D}.$
The CRM method has proved to be quite superior to classical projection methods for solving the CFP, achieving superlinear convergence under mild assumptions, while the classical methods are at most linearly convergent; see [4, 6, 5, 8, 9, 11, 10, 12, 13, 29, 47, 51, 50].
1.5. Approximate projections
CRM is superior to most known methods for the CFP, but still has a serious drawback: the computation of the projections onto the sets $C_{i}$. A variant of CRM which uses approximate projections in the product space setting was introduced in [3] and denoted as ACRM (Approximate CRM). We begin by defining separating operators for convex sets.